A mixture of CH4, C2H2 occupied a certain volume at a total pressure of 63 mm. The sample was burnt  to CO2 and H2O and the CO2 alone was collected and its pressure was found to be 69mm in the same volume at the same temperature as the original mixture. What fraction of the mixture was methane?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Nikit,

let x and y moles of methane and acetylene are present respectively..

then moles of CO2 = x+2y

x+y moles of methane and acetylene.

then x+y=63z

        x+2y=69z

then y=6z

   and x=57z

x:y = 57:6 = 19:2

therfore fraction of methane gas is 19/21